OpenStudy (anonymous):

what are the zeros of the function? f(x)=x^4-4x^3+3x^2

OpenStudy (anonymous):

x^4-4x^3+3x^2=x^2(x^2-4x+3)=x^2(x-3)(x-1) for the zeroes we have this expression equal to 0 hence the zeroes are 0,0,1,3 @RileyDuncan1999

OpenStudy (anonymous):

thak you again @Princer_Jones :)